# Context-Aware Code Generation Framework for Code Repositories: Local, Global, and Third-Party Library Awareness

### 03 Mar 2024 v3

@@ IV Experiment Setting
  • RQ2: Is Global Module Aware Knowledge effective in assisting LLMs to utilize global module knowledge during code generation, and what is the optimal “k” value for retrieving global functions?
@@ IV-A Foundation Models
We select GPT-3.5-Turbo-16k [48] as the default foundation LLM for this study due to its excellent performance in various SE tasks, along with its popularity and cost-efficiency. As for the embedding model, we use the ada-embedding-002 model [49], which has the capability of generating high-quality sentence embeddings, enabling us to perform effective similarity computation and retrieval operations. Similar to other LLM-empowered frameworks in SE domain, the more powerful model is selected, the better performance can be achieved, generally.
@@ IV-B Dataset Collection
While there exist some benchmark datasets designed for code generation, such as [50, 51, 35, 52, 16], they suffer from the following issues: 1) some of the existing datasets contain code repositories that are already present in the training corpora of popular LLMs, leading to an overestimation of the code repository-aware code generation capabilities. 2) existing benchmarks only evaluate class-level or small unit test cases, which is not suitable for our research goals. 3) Based on our literature review, none of existing code generation benchmark datasets can evaluate whether $a^{3}$ CodGen can effectively obtain, understand, and correctly reuse three distinct types of code repository information. Therefore, to address the potential risk of data leakage and accurately reflect the capability of $A^{3}$- CodGen, we created a new benchmark dataset, dubbed RepoEval. This decision stems from our intention to verify that the model truly grasp $A^{3}$ code generation after receiving information we provide.
@@ IV-B1 Code Repository Collection
Then, we create 29 function bases and 29 third-party library bases for the corresponding code repositories, following the method outlined in Section III-A. The function bases from these 29 code repositories encompass a total of 13,784 functions. Out of these, 10,077 are class member functions, leaving 3,707 as module functions. Since class member functions often involve complex relationships related to inheritance and other classes, it becomes highly labor-intensive to analyze the code invocation relationships to construct the benchmark. Therefore, our benchmark dataset focuses solely on module functions. From these 3,707 module functions, we filter out functions that had comments and were non-empty, and finally, we obtain 1,533 functions.
@@ IV-B2 Benchmark Construction
Given a function, code reuse falls into one of the three categories: 1) It could be part of locally defined functions within the local module; 2) belong to a global module; or 3) belong to a third-party library. To obtain this reuse information, we need to analyze the call relationships between functions among codes. However, existing code analysis tools like pycallgraph [53] require code compilability to accurately generate call graphs for functions. Considering that it is impractical to set up running environments for all these code repositories due to diverse requirements and potential conflicts, and given that static analysis tools like pyan [54] have relatively lower accuracy and cannot meet our needs, we resort to manual annotation to extract the call information between functions.
Manual annotation for each function is resource-intensive and time-consuming. To ensure experiment efficiency, we employ a statistical sampling method [55]. Specifically, we perform sampling on the 1,533 module functions obtained in Section IV-B1. We ensure that the estimated accuracy within the sample remains within a 0.05 error range at a 95% confidence level while ensuring functions from each code repository are included in the sample. Finally, we acquire 383 functions for subsequent experiments.
@@ IV-C Metrics
    • F1: The harmonic mean of Precision and Recall, providing an overall measure of the model’s effectiveness in reusing code, calculated as $2*(Precision*Recall)/(Precision+Recall)$.
@@
Subsequently, we enlist two developers with more than five years of experience in Python programming to conduct independent evaluations. For each generated result, they assess whether the function utilizes local functions, assigning a score of 1 if it does and 0 if it doesn’t. This score contributes to the calculation of reuse awareness. In instances of local function reuse, they identify all the reused functions, forming the basis for calculating reuse correctness. The same process applies to global functions and third-party libraries. After completing their evaluations, they discuss any discrepancies and reach a consensus. In all upcoming experiments, we employ this benchmark for evaluation and annotate the model-generated code using the same method. The Cohen’s Kappa coefficient [56] for annotated data across research questions ranges from 86% to 95%, showing high agreement among assessments.
@@ V-A The Effectiveness of Local-Aware Knowledge (RQ1)
We use vanilla ChatGPT as a baseline, which does not have any local information in its input. We define four variants of local-aware methods, as shown in Table I.
@@ V-A1 Result Analysis
Table I shows the reuse awareness results for code generation using vanilla ChatGPT and four types of Local-Aware Knowledge Configurations. Vanilla ChatGPT exhibits a precision and recall of 0 as it does not recognize any local functions. Consequently, in its generated code, new functions are often defined to complete certain logic, resulting in code redundancy.
@@ V-B The effectiveness of Global-Module-Aware Knowledge (RQ2)
Specifically, we employ No Global-Aware as the baseline and this involves providing only Local-Aware Knowledge without Global-Aware Knowledge input. Furthermore, to determine the optimal top-k value, we define four variants, namely, k=1, 5, 10, and 15. We run the total five different configurations on the benchmark to obtain code generation results. Subsequently, we assess the generated code using the same method as in section V-A, where each result is evaluated for whether it uses global functions.
@@ V-B1 Result Analysis
Table II presents the results obtained using “No Global-Aware” and four different top-k values for retrieving global functions. The No Global-Aware has both precision and recall as 0 because of lacking of global-aware information as we expected. Consequently, the generated code only utilizes local functions to accomplish tasks, and it resorts to defining new functions for specific logic if local functions are insufficient. When k=1, the model shows some awareness of global functions to a certain extent, underlined by the precision and recall as 0.384 and 0.516, respectively. At k=5, the model’s performance significantly improves, indicating that the model now has a more comprehensive understanding of global-aware information. However, when k=10, the model’s F1 score drops by 14.05%, indicating a substantial reduction in the model’s ability to reuse global functions. As k rises to 15, the effectiveness of the model decreases further. This is because the overdosed amount (between 10 to 20 functions) of functions is retrieved, introducing excessive interference and hindering accurate reusing.
@@ V-C The effectiveness of Third-Party-Library-Aware Knowledge (RQ3)
In section III-B3, we delve into Third-Party-Library Aware information, which helps the model recognize pre-installed third-party libraries in the code repository. To validate that point, we compare code generation between with and without Third-Party-Library Aware information. Only Local-Aware and Global-Aware information is fed into the model. Then, we run these two different configurations on the benchmark to obtain code generation results. Notably, we introduce the Library Coverage, the percentage of third-party libraries used in the generated code in the third-part-library base, to further elaborate the extent of reusing.
@@ V-C1 Result Analysis
Table III demonstrates that after incorporating Third-Party-Library Aware Knowledge, the model exhibits significant improvements in precision, recall, F1 score, and accuracy, with increases of 31.64%, 37.47%, 35.13%, and 16.13%, respectively. This enhancement can be attributed to the fact that, with Third-Party-Library Aware information, the model is suggested to use a specific subset of libraries, narrowing down its choice range and making it more likely to generate desired code. Our finding is further proven by that methods with Third-Party-Library Aware information outperform those without such knowledge in the Library Coverage metric by 7.43%. However, library coverage does not reach 100% due to occasional instances where the model obstinately believes that there are no viable libraries in the third-party repository to implement the code, which could be attributed to model hallucinations [57, 58, 59].
@@ V-D1 Result Analysis
Upon evaluating the awareness aspect, as shown in Table IV, $A^{3}$-CodGen demonstrates excellence in local reuse, global reuse, and third-party library reuse. This underscores the effective assistance provided by the three categories of repository-aware knowledge in enabling the model to be cognizant of repository information during code generation.
@@
Looking at the Library Coverage, $A^{3}$-CodGen achieves near-perfect performance with a score of 0.940, signifying its ability to effectively reuse pre-installed libraries and avoid compatibility issues such as “ModuleNotFoundError: No module named xxx.” Furthermore, the framework consistently generates code with an optimal number of lines (LOC), indicating its capability to produce concise and efficient code.
@@ VI Discussion
Knowledge Engineering for Code Generation. LLMs are sensitive to prompts and well-designed prompts are decisive in obtaining high-quality results [60, 61, 62]. Either unorganized or too lengthy prompts are impractical. Instead, it is advisable to extract semantic-enriched knowledge from the code repository to guide the model to directional generate code. However, how to selectively fuse the different types of knowledge presents a challenge, potentially impacting the model’s awareness of different knowledge types. In Section V-A1, we find that providing diverse information does not necessarily improve model performance. Meanwhile, in section V-D, we find that providing the model with global awareness and third-party library information unexpectedly enhances its ability to reuse native functions. This underscores the interaction of diverse code knowledge types and stresses the significance of efficiently mining and utilizing code knowledge in future studies.
Threats to Validity. The threats to validity of this paper lies in two aspects: 1) Subjective Labeling. Our study involves manual labeling, introducing subjectivity in annotations. We mitigate this by engaging two professional developers for labeling, resolving any instances with ambiguous labels to establish a reliable ground truth. The dataset is public for community validation [19]. 2) Limited Dataset Selection. Experiments use 29 meticulously selected Python code repositories, emphasizing high quality indicated by stars and forks, while avoiding data leaks. These repositories collectively represent the Python programming ecosystem. Rigorous sampling ensures a confidence level of 95% or higher, ensuring the credibility of the ground truth. In the future, we’ll maintain this dataset, providing the community with a valuable resource for evaluating repository-aware code generation.
@@ VII Related Work
Repository-Level Code Completion. Repository-level code completion is gaining traction over traditional function-level completion [63, 39, 37, 64]. Shrivastava et al. generated example-specific prompts using a prompt proposal classifier [63] and further proposed RepoFusion to incorporate relevant repository context by training code models [39]. In addition, Ding et al. [37] proposed a cross-file context finder tool, named CCFinder, which constructs a project context graph and retrieves the most relevant cross-file context based on local import statements. RepoCoder [64] retrieved similar code snippets from the repository as part of a prompt and proposes an iterative retrieval-generation paradigm. This novel strategy bridged the gap between the retrieved context and the intended target through an interactive process.
@@
Retrieval Augmented Code Generation. Retrieval Augmented Code Generation (RACG) [65, 62, 66, 67] represents an emerging and promising area of research that combines the strengths of retrieval-based methods with code generation techniques. Hayati et al. [62] introduced the RECODE model, which is the first to incorporate information retrieval techniques into the code generation task. Parvez et al. [66] proposed the REDCODER, which utilizes retrieval to enhance the input of the generation model, aiming to maximize the utilization of additional information. Drain et al. [65] retrieve template functions from GitHub and relevant answers from Stack Overflow to improve code generation. Zhou et al. [67] proposed Docprompting, which retrieve relevant APIs from documents and add them into prompts to improve code generation. Zan et al. [68] proposed CodeGenAPI, which retrieve private APIs from documents to improve code generation. the APIRetriever module of CodeGenAPI and our approach both focus on private API recommendation. However, CodeGenAPI directly retrieves APIs from documentation and cannot effectively handle high-level tasks that require combinations of APIs.
@@
Code Completion Evaluation. LLM-based code completion and program synthesis methods provide unprecedented convenience, however researcher also question the actual performance under the mirage from different perspectives [50, 51, 35, 52, 16]. Yuan et al. evaluated 10 open-source instructed LLMs on four representative code comprehension and generation tasks [50]. Liu et al. proposed EvalPlus, a code synthesis benchmarking framework, to rigorously evaluate the functional correctness of LLM-synthesized code in function level [51] . In a more challenging code generation scenario, i.e., class-level code generation, Du et al. manually constructed the first class-level code generation benchmark ClassEval of 100 class-level Python code generation tasks [35]. In addition, Ouyang et al. conducted an empirical study to demonstrate that non-determinism in code generation tasks [52]. Yu et al. proposed a benchmark named CoderEval, which can be used to assess the performance of models against pragmatic code generation beyond just generating standalone functions [16].

### 05 Mar 2024 v4

@@ IV Experiment Setting
    RQ2: Is Global Module Aware Knowledge effective in assisting LLMs to utilize global module knowledge during code generation, and what is the optimal ā€œkā€ value for retrieving global functions?
@@ IV-A Foundation Models
We select GPT-3.5-Turbo-16kĀ [48] as the default foundation LLM for this study due to its excellent performance in various SE tasks, along with its popularity and cost-efficiency. As for the embedding model, we use the ada-embedding-002 modelĀ [49], which has the capability of generating high-quality sentence embeddings, enabling us to perform effective similarity computation and retrieval operations. Similar to other LLM-empowered frameworks in SE domain, the more powerful model is selected, the better performance can be achieved, generally.
@@ IV-B Dataset Collection
While there exist some benchmark datasets designed for code generation, such asĀ [50, 51, 35, 52, 16], they suffer from the following issues: 1) some of the existing datasets contain code repositories that are already present in the training corpora of popular LLMs, leading to an overestimation of the code repository-aware code generation capabilities. 2) existing benchmarks only evaluate class-level or small unit test cases, which is not suitable for our research goals. 3) Based on our literature review, none of existing code generation benchmark datasets can evaluate whether $a^{3}$ CodGen can effectively obtain, understand, and correctly reuse three distinct types of code repository information. Therefore, to address the potential risk of data leakage and accurately reflect the capability of $A^{3}$- CodGen, we created a new benchmark dataset, dubbed RepoEval. This decision stems from our intention to verify that the model truly grasp $A^{3}$ code generation after receiving information we provide.
@@ IV-B1 Code Repository Collection
Then, we create 29 function bases and 29 third-party library bases for the corresponding code repositories, following the method outlined in SectionĀ III-A. The function bases from these 29 code repositories encompass a total of 13,784 functions. Out of these, 10,077 are class member functions, leaving 3,707 as module functions. Since class member functions often involve complex relationships related to inheritance and other classes, it becomes highly labor-intensive to analyze the code invocation relationships to construct the benchmark. Therefore, our benchmark dataset focuses solely on module functions. From these 3,707 module functions, we filter out functions that had comments and were non-empty, and finally, we obtain 1,533 functions.
@@ IV-B2 Benchmark Construction
Given a function, code reuse falls into one of the three categories: 1) It could be part of locally defined functions within the local module; 2) belong to a global module; or 3) belong to a third-party library. To obtain this reuse information, we need to analyze the call relationships between functions among codes. However, existing code analysis tools like pycallgraphĀ [53] require code compilability to accurately generate call graphs for functions. Considering that it is impractical to set up running environments for all these code repositories due to diverse requirements and potential conflicts, and given that static analysis tools like pyanĀ [54] have relatively lower accuracy and cannot meet our needs, we resort to manual annotation to extract the call information between functions.
Manual annotation for each function is resource-intensive and time-consuming. To ensure experiment efficiency, we employ a statistical sampling methodĀ [55]. Specifically, we perform sampling on the 1,533 module functions obtained in SectionĀ IV-B1. We ensure that the estimated accuracy within the sample remains within a 0.05 error range at a 95% confidence level while ensuring functions from each code repository are included in the sample. Finally, we acquire 383 functions for subsequent experiments.
@@ IV-C Metrics
[…]
@@
Subsequently, we enlist two developers with more than five years of experience in Python programming to conduct independent evaluations. For each generated result, they assess whether the function utilizes local functions, assigning a score of 1 if it does and 0 if it doesn’t. This score contributes to the calculation of reuse awareness. In instances of local function reuse, they identify all the reused functions, forming the basis for calculating reuse correctness. The same process applies to global functions and third-party libraries. After completing their evaluations, they discuss any discrepancies and reach a consensus. In all upcoming experiments, we employ this benchmark for evaluation and annotate the model-generated code using the same method. The Cohen’s Kappa coefficientĀ [56] for annotated data across research questions ranges from 86% to 95%, showing high agreement among assessments.
@@ V-A The Effectiveness of Local-Aware Knowledge (RQ1)
We use vanilla ChatGPT as a baseline, which does not have any local information in its input. We define four variants of local-aware methods, as shown in TableĀ I.
@@ V-A1 Result Analysis
TableĀ I shows the reuse awareness results for code generation using vanilla ChatGPT and four types of Local-Aware Knowledge Configurations. Vanilla ChatGPT exhibits a precision and recall of 0 as it does not recognize any local functions. Consequently, in its generated code, new functions are often defined to complete certain logic, resulting in code redundancy.
@@ V-B The effectiveness of Global-Module-Aware Knowledge (RQ2)
Specifically, we employ No Global-Aware as the baseline and this involves providing only Local-Aware Knowledge without Global-Aware Knowledge input. Furthermore, to determine the optimal top-k value, we define four variants, namely, k=1, 5, 10, and 15. We run the total five different configurations on the benchmark to obtain code generation results. Subsequently, we assess the generated code using the same method as in sectionĀ V-A, where each result is evaluated for whether it uses global functions.
@@ V-B1 Result Analysis
TableĀ II presents the results obtained using ā€œNo Global-Awareā€ and four different top-k values for retrieving global functions. The No Global-Aware has both precision and recall as 0 because of lacking of global-aware information as we expected. Consequently, the generated code only utilizes local functions to accomplish tasks, and it resorts to defining new functions for specific logic if local functions are insufficient. When k=1, the model shows some awareness of global functions to a certain extent, underlined by the precision and recall as 0.384 and 0.516, respectively. At k=5, the model’s performance significantly improves, indicating that the model now has a more comprehensive understanding of global-aware information. However, when k=10, the model’s F1 score drops by 14.05%, indicating a substantial reduction in the model’s ability to reuse global functions. As k rises to 15, the effectiveness of the model decreases further. This is because the overdosed amount (between 10 to 20 functions) of functions is retrieved, introducing excessive interference and hindering accurate reusing.
@@ V-C The effectiveness of Third-Party-Library-Aware Knowledge (RQ3)
In sectionĀ III-B3, we delve into Third-Party-Library Aware information, which helps the model recognize pre-installed third-party libraries in the code repository. To validate that point, we compare code generation between with and without Third-Party-Library Aware information. Only Local-Aware and Global-Aware information is fed into the model. Then, we run these two different configurations on the benchmark to obtain code generation results. Notably, we introduce the Library Coverage, the percentage of third-party libraries used in the generated code in the third-part-library base, to further elaborate the extent of reusing.
@@ V-C1 Result Analysis
TableĀ III demonstrates that after incorporating Third-Party-Library Aware Knowledge, the model exhibits significant improvements in precision, recall, F1 score, and accuracy, with increases of 31.64%, 37.47%, 35.13%, and 16.13%, respectively. This enhancement can be attributed to the fact that, with Third-Party-Library Aware information, the model is suggested to use a specific subset of libraries, narrowing down its choice range and making it more likely to generate desired code. Our finding is further proven by that methods with Third-Party-Library Aware information outperform those without such knowledge in the Library Coverage metric by 7.43%. However, library coverage does not reach 100% due to occasional instances where the model obstinately believes that there are no viable libraries in the third-party repository to implement the code, which could be attributed to model hallucinationsĀ [57, 58, 59].
@@ V-D1 Result Analysis
Upon evaluating the awareness aspect, as shown in TableĀ IV, $A^{3}$-CodGen demonstrates excellence in local reuse, global reuse, and third-party library reuse. This underscores the effective assistance provided by the three categories of repository-aware knowledge in enabling the model to be cognizant of repository information during code generation.
@@
Looking at the Library Coverage, $A^{3}$-CodGen achieves near-perfect performance with a score of 0.940, signifying its ability to effectively reuse pre-installed libraries and avoid compatibility issues such as ā€œModuleNotFoundError: No module named xxx.ā€ Furthermore, the framework consistently generates code with an optimal number of lines (LOC), indicating its capability to produce concise and efficient code.
@@ VI Discussion
Knowledge Engineering for Code Generation. LLMs are sensitive to prompts and well-designed prompts are decisive in obtaining high-quality resultsĀ [60, 61, 62]. Either unorganized or too lengthy prompts are impractical. Instead, it is advisable to extract semantic-enriched knowledge from the code repository to guide the model to directional generate code. However, how to selectively fuse the different types of knowledge presents a challenge, potentially impacting the model’s awareness of different knowledge types. In SectionĀ V-A1, we find that providing diverse information does not necessarily improve model performance. Meanwhile, in sectionĀ V-D, we find that providing the model with global awareness and third-party library information unexpectedly enhances its ability to reuse native functions. This underscores the interaction of diverse code knowledge types and stresses the significance of efficiently mining and utilizing code knowledge in future studies.
Threats to Validity. The threats to validity of this paper lies in two aspects: 1) Subjective Labeling. Our study involves manual labeling, introducing subjectivity in annotations. We mitigate this by engaging two professional developers for labeling, resolving any instances with ambiguous labels to establish a reliable ground truth. The dataset is public for community validationĀ [19]. 2) Limited Dataset Selection. Experiments use 29 meticulously selected Python code repositories, emphasizing high quality indicated by stars and forks, while avoiding data leaks. These repositories collectively represent the Python programming ecosystem. Rigorous sampling ensures a confidence level of 95% or higher, ensuring the credibility of the ground truth. In the future, we’ll maintain this dataset, providing the community with a valuable resource for evaluating repository-aware code generation.
@@ VII Related Work
Repository-Level Code Completion. Repository-level code completion is gaining traction over traditional function-level completionĀ [63, 39, 37, 64]. Shrivastava et al. generated example-specific prompts using a prompt proposal classifierĀ [63] and further proposed RepoFusion to incorporate relevant repository context by training code modelsĀ [39]. In addition, Ding et al.Ā [37] proposed a cross-file context finder tool, named CCFinder, which constructs a project context graph and retrieves the most relevant cross-file context based on local import statements. RepoCoderĀ [64] retrieved similar code snippets from the repository as part of a prompt and proposes an iterative retrieval-generation paradigm. This novel strategy bridged the gap between the retrieved context and the intended target through an interactive process.
@@
Retrieval Augmented Code Generation. Retrieval Augmented Code Generation (RACG)Ā [65, 62, 66, 67] represents an emerging and promising area of research that combines the strengths of retrieval-based methods with code generation techniques. Hayati et al.Ā [62] introduced the RECODE model, which is the first to incorporate information retrieval techniques into the code generation task. Parvez et al.Ā [66] proposed the REDCODER, which utilizes retrieval to enhance the input of the generation model, aiming to maximize the utilization of additional information. Drain et al.Ā [65] retrieve template functions from GitHub and relevant answers from Stack Overflow to improve code generation. Zhou et al.Ā [67] proposed Docprompting, which retrieve relevant APIs from documents and add them into prompts to improve code generation. Zan et al.Ā [68] proposed CodeGenAPI, which retrieve private APIs from documents to improve code generation. the APIRetriever module of CodeGenAPI and our approach both focus on private API recommendation. However, CodeGenAPI directly retrieves APIs from documentation and cannot effectively handle high-level tasks that require combinations of APIs.
@@
Code Completion Evaluation. LLM-based code completion and program synthesis methods provide unprecedented convenience, however researcher also question the actual performance under the mirage from different perspectivesĀ [50, 51, 35, 52, 16]. Yuan et al. evaluated 10 open-source instructed LLMs on four representative code comprehension and generation tasksĀ [50]. Liu et al. proposed EvalPlus, a code synthesis benchmarking framework, to rigorously evaluate the functional correctness of LLM-synthesized code in function levelĀ [51] . In a more challenging code generation scenario, i.e., class-level code generation, Du et al. manually constructed the first class-level code generation benchmark ClassEval of 100 class-level Python code generation tasksĀ [35]. In addition, Ouyang et al. conducted an empirical study to demonstrate that non-determinism in code generation tasksĀ [52]. Yu et al. proposed a benchmark named CoderEval, which can be used to assess the performance of models against pragmatic code generation beyond just generating standalone functionsĀ [16].